# Haoyu Zhang12  Qiaohui Chu12  Meng Liu3  Yunxiao Wang4  Bin Wen5  Fan Yang5 Tingting Gao5  Di Zhang5  Yaowei Wang12  Liqiang Nie1

### v1

@@ Abstract
AI personal assistants, deployed through robots or wearables, require embodied understanding to collaborate effectively with humans. Current Multimodal Large Language Models (MLLMs) primarily focus on third-person (exocentric) vision, overlooking the unique aspects of first-person (egocentric) videos. Additionally, high acquisition costs limit data size, impairing MLLM performance. To address these challenges, we propose learning the mapping between exocentric and egocentric domains, leveraging the extensive exocentric knowledge within existing MLLMs to enhance egocentric video understanding. To this end, we introduce Ego-ExoClip, a pre-training dataset comprising 1.1M synchronized ego-exo clip-text pairs derived from Ego-Exo4D. Our approach features a progressive training pipeline with three stages: Teacher Self-Preparation, Teacher-Student Guidance, and Student Self-Practice. Additionally, we propose an instruction-tuning data EgoIT from multiple sources to strengthen the model’s instruction-following capabilities, along with the EgoBench benchmark comprising eight different tasks for thorough evaluation. Extensive experiments across diverse egocentric tasks reveal that existing MLLMs perform inadequately in egocentric video understanding, while our model significantly outperforms these leading models [^1] https://egovisiongroup.github.io/Exo2Ego.github.io/..
1 Introduction
Figure 1: Our Exo2Ego model achieves optimal results across various egocentric video understanding tasks, with the results on EgoSchema [58], EgoTaskQA [31], and EgoMCQ [49] coming from VideoChat2 [43] and the rest from VideoLLaMA2 [11].
@@ 1 Introduction
Embodied cognition theory posits that cognitive processes are fundamentally shaped by interactions with the physical environment [34]. In this context, understanding egocentric videos that capture human active perception and fine-grained action from a first-person perspective provides a vital pathway for artificial intelligence to understand human experiences. The challenge of analyzing egocentric videos arises from the dynamic interplay between the movements of the camera-wearer (i.e., self) and the surrounding environment (i.e., other), which distinguishes them from exocentric videos observed from a third-person perspective, where this direct experience is absent. The potential applications of this technology span a wide range, including visual aids, smart glasses, and immersive experiences in virtual and augmented reality.
@@
In response to the increasing demand for effective egocentric video analysis, several research tasks have emerged, such as egocentric action recognition [40, 80, 64, 74], episodic memory [67, 68, 17, 21], human-object interaction [36, 33, 88, 3], and action anticipation [86, 70, 65]. However, the diverse architectures employed across these tasks lead to significant fragmentation in the field. With the success of existing Multimodal Large Language Models (MLLMs) [11, 42, 43], there is an emerging trend toward developing a unified egocentric MLLM architecture capable of addressing multiple tasks within the egocentric domain. A significant obstacle in this direction is the limited availability of egocentric video data, which is crucial for effectively training large-capacity models [80]. Consequently, an increasing number of methods [45, 89, 52, 19] are now retrieving suitable exocentric videos from the predefined set to aid in model training for downstream tasks. While these approaches exhibit empirical advantages in transferring knowledge from exocentric to egocentric representations, they incur additional retrieval time and suffer from temporal misalignment, leading to instability in model performance.
@@
Figure 2: Understanding how humans map observed activities to their own behaviors enhances embodied cognition across various downstream tasks by transferring insights from existing Internet (i.e., exocentric) data [3].
@@
Drawing inspiration from cognitive science [22, 73], which demonstrates that children learn by observing others’ behavior (i.e., an exocentric view) and mapping these observations onto their own experiences (i.e., an egocentric perspective), we conceptualize the exocentric video observer as the “teacher” and the egocentric video interpreter as the “student”. Our objective is to establish a robust mapping pattern between the student and teacher, as illustrated in Figure 2, offering two key advantages: 1) Weak Dependence. Leveraging the learned mappings, the model can flexibly tackle downstream tasks without relying on cross-domain data, thereby mitigating additional biases. 2) Strong Generalization. By emulating human learning, these mapping techniques significantly enhance the model’s generalization capabilities, reducing the need for extensive egocentric training data in MLLMs.
To further this approach, we propose a semantic mapping framework grounded in cross-view behavior invariance. Our framework integrates an exocentric-to-egocentric migration module with an egocentric self-consistency mechanism. By embedding this framework within MLLMs, we design a progressive training pipeline comprising three stages: Teacher Self-Preparation, Teacher-Student Guidance, and Student Self-Practice. This pipeline enables effective teacher-student mappings that transfer exocentric knowledge to improve egocentric vision comprehension, thereby enhancing both the effectiveness and generalization of the model in downstream tasks.
Furthermore, to satisfy the demand for synchronized cross-domain data, we introduce the Ego-ExoClip dataset, an advanced first- and third-person video-text dataset containing 1.1M egocentric-exocentric clip-text pairs. Ego-ExoClip is, to the best of our knowledge, the largest and most behaviorally diverse clip-text collection available in this domain.
To further improve the instruction-following capabilities of MLLMs, we construct an enriched instruction-tuning dataset, EgoIT, sourced from multiple egocentric contexts. In addition, we propose the EgoBench benchmark, which encompasses eight distinct tasks across more than ten settings to evaluate the performance of existing MLLMs. As shown in Figure 1, experimental results on EgoBench reveal the suboptimal performance of current MLLMs (e.g., VideoLLaMA2 [11] and VideoChat2 [43]), whereas our framework achieves superior results. This validation underscores the effectiveness of our approach in advancing egocentric video understanding and promotes further research in embodied cognition.
2 Related Work
Figure 3: Examples from the proposed pre-training Ego-ExoClip dataset, instruction-tuning data EgoIT, and benchmark EgoBench.
@@ 2 Related Work
Egocentric Video Understanding. Egocentric videos offer a distinctive perspective for active engagement with the physical world, capturing human interactions from a first-person viewpoint. A variety of datasets have been developed to support research in this field, including EPIC-KITCHENS [14], which focuses on kitchen activities; Charades-Ego [75], which highlights diverse everyday tasks; and Ego4D [24], which provides a globally diverse collection of egocentric videos. These datasets have driven the growth of multiple research areas, such as video-text pre-training [49, 63], video question answering [31, 18, 96, 95], human-object interaction [36, 33, 88, 3], action recognition [40, 80, 64, 74, 47], action anticipation [86, 70, 65, 72], episodic memory [67, 68, 17, 21], pose perception [2, 60, 35, 51], and visual queries with 2D/3D localization [32, 55]. Despite these advancements, the high cost of data collection remains a major limitation, constraining dataset scale and, in turn, the effectiveness of the models trained on them. In response, recent work [45, 82, 89, 78] has begun exploring the use of unpaired exocentric videos to enhance egocentric video comprehension. However, leveraging unpaired exocentric data introduces inherent biases that can negatively affect model stability and accuracy. To address this issue, we introduce a pairwise egocentric-exocentric dataset with diverse activities and leverage the exocentric knowledge embedded in MLLMs to improve egocentric video understanding.
@@
Multimodal Large Language Model. Large Language Models (LLMs) [7, 77, 1, 12], trained on vast datasets from the Internet, have demonstrated remarkable proficiency across a wide range of language tasks through text generation. Building upon the success of LLMs and the promise of multimodal data, there is growing scholarly interest in exploring MLLMs. Early MLLMs focus on integrating static images with text, with landmark work such as Flamingo [4], PaLM-E [20], and BLIP-2 [41], significantly expanding the capabilities of LLMs for multimodal tasks. Recent open-source MLLMs, including LLaVA [50], MiniGPT-4 [101], and InstructBLIP [13], further advance visual understanding by introducing visual instruction-tuning data. Beyond static images, recent studies have started to explore the incorporation of video data into LLMs, unlocking new potential for video comprehension tasks. Innovations such as LLaMA-VID [46], Video-LLaVA [48], and VideoLLaMA2 [11] aim to improve the alignment of video data with LLMs. Additionally, methods like VideoChat [42], VideoChatGPT [54], Valley [53], and VideoChat2 [43] utilize ChatGPT [^2] https://chatgpt.com/. to generate video instruction-tuning data, enhancing instruction-following capabilities. Despite these advancements, existing MLLMs exhibit limited performance on egocentric video understanding tasks [80]. Therefore, our objective is to unify various downstream tasks and significantly enhance the capabilities of MLLMs in the embodied cognition domain.
3 Data Construction
@@
Figure 4: Illustration of our progressive training process, including the stages of Teacher Self-Preparation, Teacher-Student Guidance, and Student Self-Practice. “VTG”, “KL”, and “CCL” denote vision-grounded text generation, Kullback-Leibler divergence, and cycle consistency loss, respectively. $F$ is a mapping function.
3.1 Ego-ExoClip
The Ego-ExoClip dataset is derived from Ego-Exo4D [25], which comprises 5,035 grouped videos ranging in length from 1 to 42 minutes. Each group contains one egocentric video paired with 4–5 simultaneously captured exocentric videos. Most of these videos include dense timestamp-level narrations provided by two annotators, detailing the camera wearer’s activities and interactions. For example, the narration “C pours some water from the cup into the skillet with her right hand” corresponds to an event occurring at 622.47 seconds, where “C” refers to the camera-wearer.
To ensure data quality and consistency, we filter out videos without narrations, those lacking UID mappings [^3] UID mapping associate narrations and corresponding video files., and those designated for validation or testing in benchmark challenges. This process yields a refined dataset of 2,925 video groups, totaling 15,478 videos, all validated to guarantee lossless quality. For incorporating narration diversity, we retain textual annotations from both narrators, resulting in 623.6 hours of videos and 261.3K narrations.
@@ 3.1 Ego-ExoClip
where $\beta_{i}$ is a parameter representing the average temporal distance between consecutive narrations, calculated as $\sum_{j=0}^{n-1}(t_{j+1}-t_{j})/n$. This value is calculated for each video individually, with $\alpha$ serving as a scale factor derived from the average of all $\beta_{i}$ across the entire Ego-ExoClip dataset ($\alpha$ = 1.92 seconds). To ensure that a description does not span multiple clips, we utilize the preceding and following timestamps as boundary markers. A data sample from Ego-ExoClip is illustrated in Figure 3, with further details and statistical analyses available in Supp. A.
3.2 EgoIT
To enhance the instruction-following capabilities of MLLMs for downstream egocentric tasks, we introduce the EgoIT dataset, which comprises approximately 600K samples sourced from multiple distinct domains, as detailed in Table 1. All data samples are standardized into a uniform format, as shown in Figure 3. Each sample consists of four main components: the video path, task instruction, question, and answer. To ensure diversity in instructions, we integrate the “dataset description”, “task description”, and “instruction example” information into a prompt template, using ChatGPT [^4] https://openai.com/index/hello-gpt-4o/. to generate ten distinct instructions per dataset. Further details are provided in Supp. B.
Our instruction-tuning EgoIT dataset can be broadly categorized into three areas: 1) Action Recognition, which focuses on enhancing the identification of hand-object interactions. We utilize the EGTEA [44] and Something-Something-V2 [23] datasets within the egocentric setting. 2) Question Answering, which aims to improve the ability to answer questions based on video content. We include the EgoTimeQA [18] and OpenEQA [56] datasets, encompassing both open-ended and close-ended questions. And 3) Captioning, which pays attention to enhancing basic visual description capabilities, utilizing the EgoExoLearn [29] dataset for this purpose.
4 Method
@@
Initialization Panda-70M [9], WebVid-10M [5], VIDAL-10M [100], InternVid-10M [83] 103M 3.8M
CC-3M [8], DCI-7.8K [79], EgoClip-3.8M [49]
@@
Stage 3 OpenEQA-1.6K [57], EgoTimeQA-303K [18], Something-Something-V2-221K [59] - 586K
EGTEA-15K [44], EgoExoLearn-45K [29]
@@
Table 1: Overview of the datasets utilized in the various training stages, detailing their sources and scales.
4.1 Initialization
@@
The goal of the initialization is to equip both the exocentric and egocentric visual encoders with foundational capabilities to process their respective data types. We tackle this by training the encoders separately using data from both domains. For the exocentric visual encoder, we follow the approach of VideoLLaMA2 [11] and utilize a large-scale, web-crawled dataset comprising image-text and video-text pairs from various publicly accessible databases, as summarized in Table 1. For the egocentric visual encoder, we employ the EgoClip [49] dataset, which includes a wide range of first-person daily activities. During this training phase, the parameters of LLM are frozen, and we optimize only the two visual encoders. The trained parameters of visual encoders then serve as initialization weights for the subsequent three-stage process.
4.2 Stage 1: Teacher Self-Preparation
@@
After initialization, the exocentric visual encoder (the teacher) must adapt to the Ego-ExoClip dataset, despite already being proficient at understanding exocentric videos. To achieve this, we implement a self-preparation step to ensure that the exocentric encoder is adequately prepared to transfer its knowledge to the egocentric branch. As shown in Figure 4, we freeze the LLM parameters and fine-tune the exocentric visual encoder using exocentric clip-text data from our proposed Ego-ExoClip dataset. For supervision, we employ Vision-grounded Text Generation (VTG) loss to ensure data accuracy.
4.3 Stage 2: Teacher-Student Guidance
@@
In the second stage, we aim to establish a robust correlation between the exocentric and egocentric data, effectively mapping the student to the teacher. This mapping is essential, as exocentric knowledge can significantly enhance the comprehension of egocentric videos, given that human behaviors are largely perspective-invariant and independent of camera viewpoint. For this stage, we continue to adopt Ego-ExoClip as input data. With the exocentric visual encoder sufficiently trained, we freeze it to maintain its integrity. We also freeze the LLM while fine-tuning the egocentric visual encoder and the mapping function $F$. The mapping $F:X\rightarrow Y$ is inherently under-constrained, prompting us to couple it with an inverse mapping $G:Y\rightarrow X$, where $X$ and $Y$ denote the egocentric and exocentric domains, respectively. Both mappings are treated as bijections, and we enforce this structural assumption by training them simultaneously while applying a Cycle Consistency Loss (CCL), defined as:
@@
where $x\in X$, $y\in Y$, and $\|\|_{1}$ denotes the L1 norm. The CCL encourages both forward cycle consistency, i.e., $x\rightarrow F(x)\rightarrow G(F(x))\approx x$, and backward cycle consistency, i.e., $y\rightarrow G(y)\rightarrow F(G(y))\approx y$. Additionally, we introduce the Kullback–Leibler (KL) divergence [97] between the exocentric sample $y$ and the estimated exocentric sample $\hat{y}=F(x)$, aligning real and estimated exocentric feature distributions. The concatenated guidance from the teacher and the student’s own information is then fed into the LLM, which is updated using VTG loss.
4.4 Stage 3: Student Self-Practice

### v2

# Exo2Ego: Exocentric Knowledge Guided MLLM for Egocentric Video Understanding

@@ Abstract
AI personal assistants, deployed through robots or wearables, require embodied understanding to collaborate effectively with humans. However, current Multimodal Large Language Models (MLLMs) primarily focus on third-person (exocentric) vision, overlooking the unique challenges of first-person (egocentric) videos. Additionally, high acquisition costs limit data size, impairing MLLM performance. To address these challenges, we propose learning the mapping between exocentric and egocentric domains, leveraging the extensive exocentric knowledge within existing MLLMs to enhance egocentric video understanding. To this end, we introduce Ego-ExoClip, a pre-training dataset comprising 1.1M synchronized ego-exo clip-text pairs derived from Ego-Exo4D, together with the instruction-tuning dataset EgoIT, which is collected from multiple sources to enhance the model’s instruction-following capabilities. Building upon the datasets, we propose a migration strategy and further design a progressive mapping learning pipeline with three stages: Demonstrator Self-Preparation, Demonstrator-Learner Guidance, and Learner Self-Practice. Extensive experiments across diverse egocentric tasks reveal that existing MLLMs perform inadequately in egocentric video understanding, while our model significantly outperforms these leading models.
Codehttps://reurl.cc/Ebpyrm
Introduction
Figure 1: Our Exo2Ego model achieves optimal results across various egocentric video understanding tasks, with the detailed results presented in Table 2.
@@ Introduction
Embodied cognition theory posits that cognitive processes are fundamentally shaped by interactions with the physical environment (10.3389/fpsyg.2015.00875). In this context, understanding egocentric videos that capture human active perception and fine-grained action from a first-person perspective provides a vital pathway for artificial intelligence to understand human experiences. The challenge of analyzing egocentric videos arises from the dynamic interplay between the movements of the camera-wearer (i.e., self) and the surrounding environment (i.e., other), which distinguishes them from exocentric videos observed from a third-person perspective, where this direct experience is absent. The potential applications of this technology span a wide range, including visual aids, smart glasses, and immersive experiences in virtual and augmented reality.
@@
In response to the increasing demand for effective egocentric video analysis, several research tasks have emerged, such as egocentric action recognition (shiota2024egocentric; pmlr-v235-zhang24aj), episodic memory (feng2024objectnlq; wang2022siamese; Feng_2025_CVPR), human-object interaction (jiang2023full), and action anticipation (chu2025intention; chu2025technical). However, the diverse architectures employed across these tasks lead to significant fragmentation in the field. With the success of existing Multimodal Large Language Models (MLLMs) (wang2025time; zhang2025spatial), there is an emerging trend toward developing a unified egocentric MLLM architecture capable of addressing multiple tasks within the egocentric domain. A significant obstacle in this direction is the limited availability of egocentric video data, which is crucial for effectively training large-capacity models (wang2023ego). Consequently, an increasing number of methods (li2021ego; xue2023learning; luo2025put; dou2024unlocking) are now retrieving suitable exocentric videos from the predefined set to aid in model training for downstream tasks. While these approaches exhibit empirical advantages in transferring knowledge from exocentric to egocentric representations, they incur additional retrieval time and suffer from misalignment, leading to instability in model performance.
@@
Figure 2: Understanding how humans map observed activities to their own behaviors enhances embodied cognition across various downstream tasks by transferring insights from existing Internet (i.e., exocentric) data.
@@
Drawing inspiration from cognitive science (shi2024cognition; zhang2023attribute), which demonstrates that children learn by observing others’ behavior (i.e., an exocentric view) and mapping these observations onto their own experiences (i.e., an egocentric perspective), we conceptualize the exocentric video observer as the “demonstrator” and the egocentric video interpreter as the “learner”. Our objective is to establish a robust mapping pattern between the two, as illustrated in Figure 2, offering two key advantages: 1) Weak Dependence. Leveraging the learned mappings, the model can flexibly tackle downstream tasks without relying on cross-domain data, thereby mitigating additional biases. 2) Strong Generalization. By emulating human learning, these mapping techniques significantly enhance the model’s generalization capabilities, reducing the need for extensive egocentric training data in MLLMs.
To further this approach, we propose an exocentric-to-egocentric migration strategy with an egocentric self-consistency mechanism grounded in cross-view behavior invariance. By embedding this strategy within MLLMs, we design a mapping learning pipeline comprising three stages: Demonstrator Self-Preparation, Demonstrator-Learner Guidance, and Learner Self-Practice. By transferring exocentric knowledge, this pipeline boosts egocentric understanding and improves downstream performance. Furthermore, to satisfy the demand for synchronized cross-domain data, we introduce the Ego-ExoClip dataset, an advanced first- and third-person video-text dataset containing 1.1M egocentric-exocentric clip-text pairs. And we also construct an enriched instruction-tuning dataset, EgoIT, sourced from multiple egocentric contexts to improve the instruction-following capabilities. As shown in Figure 1, abundant results demonstrate that our framework achieves superior results, highlighting its effectiveness in advancing egocentric video understanding.
Contributions: 1) Based on behavior invariance, we propose a knowledge transfer strategy and integrate it into the training paradigm of MLLMs, resulting in a three-stage mapping learning process. 2) To support training, we construct a synchronized exo-ego video-text dataset, Ego-ExoClip, as well as an instruction-tuning dataset, EgoIT, collected from multiple sources. 3) Extensive experimental results on various benchmarks validate the effectiveness of our proposed datasets and methods, and further promote research in embodied cognition.
Related Work
@@
Egocentric Video Understanding. Egocentric videos offer a distinctive perspective for active engagement with the physical world, capturing human interactions from a first-person viewpoint. A variety of datasets have been developed to support research in this field, including EPIC-KITCHENS (damen2020epic), which focuses on kitchen activities; Charades-Ego (sigurdsson2018charades), which highlights diverse everyday tasks; and Ego4D (grauman2022ego4d), which provides a globally diverse collection of egocentric videos. These datasets have driven the growth of multiple research areas, such as video-text pre-training (lin2022egocentric; pramanick2023egovlpv2), video question answering (di2024grounded; zhang2024hcqa; zhang2021multimodal), human-object interaction (akiva2023self), action anticipation (ragusa2023stillfast; shen2024progress), episodic memory (ramakrishnan2023naq; feng2024objectnlq), and pose perception (millerdurai2024eventego3d). Despite these advancements, the high cost of data collection remains a major limitation, constraining dataset scale and, in turn, the effectiveness of the models trained on them. In response, recent work (wang2023learning; xue2023learning; truong2024cross) has begun exploring the use of unpaired exocentric videos to enhance egocentric video comprehension. However, leveraging unpaired exocentric data introduces inherent biases that can negatively affect model stability and accuracy. To address this issue, we introduce a pairwise egocentric-exocentric dataset with diverse activities and leverage the exocentric knowledge embedded in MLLMs to improve egocentric video understanding.
@@
Multimodal Large Language Model. Large Language Models (LLMs) (touvron2023llama; DBLP:journals/pami/WenSYWGN24), trained on vast datasets from the Internet, have demonstrated remarkable proficiency across a wide range of language tasks through text generation. Building upon the success of LLMs and the promise of multimodal data, there is growing scholarly interest in exploring MLLMs. Early MLLMs focus on integrating static images with text, with landmark work such as Flamingo (alayrac2022flamingo) and BLIP-2 (li2023blip), significantly expanding the capabilities of LLMs for multimodal tasks. Recent open-source MLLMs, including LLaVA (liu2024visual) and InstructBLIP (NEURIPS2023_9a6a435e), further advance visual understanding by introducing visual instruction-tuning data. Beyond static images, recent studies have started to explore the incorporation of video data into LLMs, unlocking new potential for video comprehension tasks. Innovations such as LLaMA-VID (li2023llama), Video-LLaVA (lin2023video), and VideoLLaMA2 (cheng2024videollama) aim to improve the alignment of video data with LLMs. Additionally, methods like VideoChatGPT (maaz2023video) and VideoChat2 (li2024mvbench) utilize ChatGPT to generate video instruction-tuning data, enhancing instruction-following capabilities. Despite these advancements, existing MLLMs exhibit limited performance on egocentric video understanding tasks (wang2023ego). Therefore, our objective is to unify various downstream tasks and significantly enhance the capabilities of MLLMs in the embodied cognition domain.
Dataset Construction
@@
Figure 3: Illustration of the construction of Ego-ExoClip.
Ego-ExoClip
The Ego-ExoClip dataset is derived from Ego-Exo4D (grauman2024ego), which comprises 5,035 grouped videos ranging in length from 1 to 42 minutes. Each group contains one egocentric video paired with 4–5 simultaneously captured exocentric videos. Most of these videos include dense timestamp-level narrations provided by two annotators, detailing the camera wearer’s activities and interactions. For example, as shown in Figure 3, the narration “C cracks the egg open in the bowl on the countertop with both hands” corresponds to an event occurring at 622.47 seconds, where “C” refers to the camera-wearer.
To ensure data quality and consistency, we filter out videos without narrations, those lacking UID mappings [^1] UID mapping associate narrations and corresponding videos., and those designated for validation or testing in benchmark challenges. This process yields a refined dataset of 2,925 video groups, totaling 15,478 videos, all validated to guarantee lossless quality. For incorporating narration diversity, we retain textual annotations from both narrators, resulting in 623.6 hours of videos and 261.3K narrations.
@@
where $\beta_{i}$ is a parameter representing the average temporal distance between consecutive narrations, calculated as $\sum_{j=0}^{n-1}(t_{j+1}-t_{j})/n$. This value is calculated for each video individually, with $\alpha$ serving as a scale factor derived from the average of all $\beta_{i}$ across the entire Ego-ExoClip dataset ($\alpha$ = 1.92 seconds). To ensure that a description does not span multiple clips, we utilize the preceding and following timestamps as boundary markers.
Figure 4: Illustration of our progressive training process, including the stages of Demonstrator Self-Preparation, Demonstrator-Learner Guidance, and Learner Self-Practice. “VTG”, “KL”, and “CCL” denote vision-grounded text generation, Kullback-Leibler divergence, and cycle consistency loss, respectively. $F$ is a mapping function.
EgoIT
To enhance the instruction-following capabilities of MLLMs for downstream egocentric tasks, we introduce the EgoIT dataset, which comprises approximately 600K samples sourced from multiple distinct domains. All data samples are standardized into a uniform format. Each sample consists of four main components: the video path, task instruction, question, and answer. To ensure diversity in instructions, we integrate the “dataset description”, “task description”, and “instruction example” information into a prompt template, using GPT-4o [^2] https://openai.com/index/hello-gpt-4o/. to generate ten distinct instructions per dataset.
Our instruction-tuning EgoIT dataset can be broadly categorized into three areas: 1) Action Recognition, which focuses on enhancing the identification of hand-object interactions. We utilize the EGTEA (li2021eye) and Something-Something-V2 (goyal2017something) datasets within the egocentric setting. 2) Question Answering, which aims to improve the ability to answer questions based on video content. We include the EgoTimeQA (di2024grounded) and OpenEQA (Majumdar_2024_CVPR) datasets, encompassing both open-ended and close-ended questions. And 3) Captioning, which pays attention to enhancing basic visual description capabilities, utilizing the EgoExoLearn (huang2024egoexolearn) dataset for this purpose.
Method
@@
Initialization Panda-70M, WebVid-10M, VIDAL-10M, InternVid-10M 103M 3.8M
CC-3M, DCI-7.8K, EgoClip-3.8M
@@
Stage 3 OpenEQA-1.6K, EgoTimeQA-303K, Something-Something-V2-221K - 586K
EGTEA-15K, EgoExoLearn-45K
@@
Table 1: Overview of the datasets utilized in the multiple training stages, detailing their sources and scales.
Method ES QAEgo4D EgoTaskQA CE EK EP VLN-QA EgoMCQ
Acc. Acc. Acc./Score Acc. Acc. mAP mAP Acc. Acc. Acc. Acc.
Specific Egocentric Methods
EgoVLP - 22.6 17.3/1.9 42.5 38.7 25.0 26.0 - - - -
EgoVLPv2 - 25.8 19.7/2.1 46.3 42.3 30.7 29.1 15.3 10.4 72.3 30.6
MFAS - 37.1 23.2/2.2 48.7 45.4 37.4 30.3 17.2 9.5 77.5 34.2
GroundVQA - 50.2 29.0/2.6 - - 42.5 - 22.7 17.4 80.3 37.1
LaViLa++ - - - - - 40.8 - 20.9 11.0 75.4 33.7
Zero-shot Closed-source MLLMs
Gemini 1.5-Pro 71.2 - - - - - - - - - -
GPT-4o 72.2 60.3 29.1/3.0 43.2 47.5 59.4 45.3 36.8 34.0 86.6 38.8
Zero-shot Open-source MLLMs
LLaMA-VID 38.5 38.6 23.5/2.0 29.0 31.6 31.0 46.1 27.2 32.3 56.5 23.5
Video-LLaVA 38.4 35.2 28.7/2.7 29.8 30.6 43.5 45.2 27.5 34.0 60.9 25.1
VideoLLaMA2 51.7 48.5 25.3/2.6 42.1 42.6 68.1 47.8 29.3 37.5 82.1 32.0
VideoChat2 54.4 45.4 21.9/1.7 42.3 47.4 57.0 42.6 26.1 35.0 85.1 37.6
Qwen2.5-VL 57.6 55.3 26.4/2.8 42.6 45.8 64.2 46.4 32.5 34.3 85.3 38.2
Exo2Ego 61.3 62.1 28.3/2.7 44.7 50.3 70.9 49.7 42.7 44.5 88.4 41.2
Table 2: Performance comparison with recent state-of-the-art methods methods on eight benchmarks, where ES, CE, EK, and EP refer to EgoSchema, Charades-Ego, EPIC-KITCHENS-100, and EgoPlan. The QAEgo4D dataset is categorized into close and open question answering, EgoTaskQA includes direct and indirect subsets, and EgoMCQ contains inter and intra settings. The best results among open-source MLLMs are highlighted in bold.
Initialization
@@
The goal of the initialization is to equip both the exocentric and egocentric visual encoders with foundational capabilities to process their respective data types. We tackle this by training the encoders separately using data from both domains. For the exocentric visual encoder, we utilize a large-scale, web-crawled dataset comprising image-text and video-text pairs from various publicly accessible databases, as summarized in Table 1. For the egocentric visual encoder, we employ the EgoClip (lin2022egocentric) dataset, which includes a wide range of first-person daily activities. During this training phase, the parameters of LLM are frozen, and we optimize only the two visual encoders. The trained parameters of visual encoders then serve as initialization weights for the subsequent three-stage process.
Stage 1: Demonstrator Self-Preparation
@@
After initialization, the exocentric visual encoder (the demonstrator) must adapt to the Ego-ExoClip dataset, despite already being proficient at understanding exocentric videos. To achieve this, we implement a self-preparation step to ensure that the exocentric encoder is adequately prepared to transfer its knowledge to the egocentric branch. As shown in Figure 4, we freeze the LLM parameters and fine-tune the exocentric visual encoder using exocentric clip-text data from our proposed Ego-ExoClip dataset. For supervision, we employ Vision-grounded Text Generation (VTG) loss to ensure data accuracy.
Stage 2: Demonstrator-Learner Guidance
@@
In the second stage, we aim to establish a robust correlation between the exocentric and egocentric data, effectively mapping the learner to the demonstrator. This mapping is essential, as exocentric knowledge can significantly enhance the comprehension of egocentric videos, given that human behaviors are largely perspective-invariant and independent of camera viewpoint. For this stage, we continue to adopt Ego-ExoClip as input data. With the exocentric visual encoder sufficiently trained, we freeze it to maintain its integrity. We also freeze the LLM while fine-tuning the egocentric visual encoder and the mapping functions $F$ and $G$. The mapping $F:X\rightarrow Y$ is inherently under-constrained, prompting us to couple it with an inverse mapping $G:Y\rightarrow X$, where $X$ and $Y$ denote the egocentric and exocentric domains, respectively. Both mappings are treated as bijections, and we enforce this structural assumption by training them simultaneously while applying a Cycle Consistency Loss (CCL), defined as:
@@
where $x\in X$, $y\in Y$, and $\|\|_{1}$ denotes the L1 norm. The CCL encourages both forward cycle consistency, i.e., $x\rightarrow F(x)\rightarrow G(F(x))\approx x$, and backward cycle consistency, i.e., $y\rightarrow G(y)\rightarrow F(G(y))\approx y$. Additionally, we introduce the Kullback–Leibler (KL) divergence between the exocentric sample $y$ and the estimated exocentric sample $\hat{y}=F(x)$, aligning real and estimated exocentric feature distributions. The concatenated guidance from the demonstrator and the learner’s own information is then fed into the LLM.
Stage 3: Learner Self-Practice